# The 2D+ Dynamic Articulatory Model DYNARTmo: Tongue-Palate Contact Area Estimation

### v3

@@ 1. Introduction
Articulatory models describe the stage of speech where neuromuscular activation patterns lead to movement patterns of speech articulators (i.e., lips, tongue, velum, lower jaw) with the goal of generating audible speech signals as part of speech communication processes. The motivation for developing articulatory models is manifold. (i) Models unfold the process of speech articulation and acoustic signal generation and thus help to understand the basic mechanisms of articulation and coarticulation. (ii) If coupled with an articulatory-acoustic module, an articulatory model provides insight into the mechanisms of acoustic signal generation at the vocal folds and within the vocal tract. (iii) Articulatory models may serve in the future as high-quality speech synthesizers, capable of generating different voices (i.e., different speakers), different emotional connotations, and more—beyond merely transmitting a spoken message. However, corpus-based synthesis systems currently produce the most natural speech quality, and many research questions remain to be addressed before high-quality articulatory synthesizers can be realized (Campbell,, 2005; Kröger,, 2022).
@@
At present, articulatory models are mainly used as research tools to study articulation and coarticulation or, when combined with articulatory-acoustic modules, to investigate aspects of acoustic signal generation, including sublaryngeal, laryngeal, and supralaryngeal aerodynamics (Birkholz,, 2013; Kröger,, 2022; Fan et al.,, 2024). Other models are used in speech therapy to visually present articulator positions ((Kröger et al.,, 2005)) or articulatory contact information (e.g., tongue–palate or labial contact), offering somatosensory feedback to help improve our imagination of articulation. Examples include electropalatography (EPG), which provides real‑time visualization of tongue–palate contact during speech (e.g. (Hardcastle et al.,, 1991; Nordberg et al.,, 2011)), and ultrasound visual biofeedback (U‑VBF), displaying dynamic tongue surface movements in therapy for speech sound disorders ((Sugden et al.,, 2019; Preston et al.,, 2017)). Moreover, articulatory models can be used to simulate speech learning or relearning processes in therapeutic contexts, or as part of second and third language learning. In that case, articulatory models need to be integrated into large-scale neurobiologically motivated neural models of language processing (Kröger,, 2023).
@@
Many aspects of speech articulation can be represented in the midsagittal plane. Therefore, early articulatory models were typically two-dimensional (Henke,, 1966; Mermelstein,, 1973; Coker,, 1976; Maeda,, 1979; Heike,, 1979). These models primarily aimed at generating high-quality synthetic speech. However, by the end of the 20th century, corpus-based synthesizers such as the Klatt synthesizer (Klatt,, 1980) and newer systems like corpus-bases speech synthesizers (Campbell,, 2005) had outperformed articulatory models in speech quality. Consequently, articulatory models—including articulatory-acoustic variants—became more focused on their use as research tools, particularly in studying speech acquisition, articulation, and coarticulation across languages and speaker groups. As a result, increasingly detailed models emerged that incorporate 3D spatial information, neuromuscular control schemes, and biomechanical modeling of muscles and articulatory tissues (Kröger,, 2022).
@@
The model described in this paper, DYNARTmo (DYNamic ARTiculatory model), is a further development of the UK-DYNAMO (Universität Köln-DYNamic Articulatory MOdel) originally introduced by Heike (Heike,, 1979). The UK-DYNAMO was inspired by Henke’s approach (Henke,, 1966) and later used in a computer-based tool for speech training (Heike,, 1989). Like many models of its time, UK-DYNAMO was a 2D system based on midsagittal contours of articulator surfaces, specifically German vowel and consonant contours from Wängler (Wängler,, 1958). It employed temporal interpolation and a basic form of coarticulation as suggested by Henke.
@@
DYNARTmo was initially introduced by Kröger (Kröger,, 1992), who incorporated the concept of articulatory underspecification to derive minimal rules for articulation and coarticulation. This approach was elaborated through two control concepts: a gestural control scheme (Kröger,, 1993; Kröger et al.,, 1995) and a segmental approach (Kröger,, 2001). Starting in the 2000s, DYNARTmo was updated with geometries derived from MRI data (Kröger et al.,, 2000, 2004), offering a more realistic articulatory basis. Later, it was embedded into a large-scale neurobiological speech production model (Kröger et al.,, 2014), incorporating concepts such as primary and secondary articulators, muscle bundles, and neuromuscular activation patterns (Kröger and Bekolay,, 2022).
@@ 2. DYNARTmo: Static Aspects
Table 1: Continuously scaled control parameters used in DYNARTmo, including affected articulators and their functional roles. Vocalic parameters define global vocal tract shaping; consonantal parameters govern localized constrictions. Additional parameters control velopharyngeal aperture, glottal configuration, and subglottal pressure. (Speech sounds are transcribed using SAMPA; see (Wells,, 1997))
@@
In addition to the ten continuously scaled control parameters (see Table 1), the model includes six discrete parameters that define the place and manner of consonantal constrictions (see Table 2). The value range of the continuous parameters varies depending on their functional category: for vocalic parameters, values range from 1-1 to $1$, where 0 indicates the neutral articulator position. The same 1-1 to $1$ scale is used for velopharyngeal and glottal aperture, with $1$ indicating full opening, 0 indicating closure, and 1-1 representing tight closure. For parameters governing consonantal constrictions (i.e., labial, apical, and dorsal), as well as for vocal fold tension and lung pressure, the scale ranges from 0 to $1$. A value of 0 indicates no active constriction (i.e., the articulator remains in its current vocalic position), while a value of $1$ represents full articulatory closure or maximum effort.
@@
tongue tip manner full (vocalic), near (central groove), lateral (lowering)
@@
The set of articulatory control parameters does not directly determine the positions of the model articulators. Rather, these parameters serve a functional role: they specify consonantal articulation in terms of the degree and location of vocal tract constrictions, and vocalic articulation in terms of canonical phonetic-linguistic categories such as high–low, front–back, and spread–rounded (cf.(IPA,, 1999): International Phonetic Association).
The actual specification of articulator positions is carried out in a subsequent computational step, based on the values of all control parameters at a given time point. This step involves computing the vertical and horizontal positions of the tongue tip, tongue dorsum, and lips, as well as the vertical positions of the velum, lower jaw, and larynx (see Fig. 1). Conceptually, this transformation can be understood as the mapping of high-level vocal tract parameters—such as constriction degree and location for consonants, or vocal tract tube shape for vowels—onto concrete articulator geometries, akin to the task dynamics approach proposed by (Saltzman and Munhall,, 1989).
@@ 3. Integration of DYNARTmo into a Web App for Speech Therapy and Language Learning
In addition to displaying the form and displacement of articulators from their rest positions in a sagittal plane view (see Fig. 1), the SpeechArticulationTrainer web application also provides: (i) a glottal view (transverse plane, superior perspective), and (ii) a palatal view (transverse plane, inferior perspective), as illustrated in Figure 2.
@@ 4. Discussion
One central goal of this study is to propose a simplified articulatory model that operates without the full task-dynamics formalism, while still reflecting key distinctions commonly found in more complex frameworks such as the Task Dynamics Approach ((Saltzman and Munhall,, 1989)). Notably, our model implicitly includes a distinction between articulator space and tract variable space. However, this distinction is not strictly separated in our formulation. For instance, articulatory parameters such as tongue dorsum height or velum position are directly mapped to constriction properties, thereby bypassing an explicit intermediate representation of tract variables. This deliberate simplification serves to make the model computationally tractable and visually transparent, especially for educational or diagnostic applications. Nonetheless, the blending of articulator-based and functionally-defined parameters might lead to a certain conceptual ”blurring” that deserves clarification. Future refinements of the model may aim to more clearly disentangle these parameter spaces, or at least highlight their interplay more explicitly, in order to facilitate theoretical comparisons with established models of speech production.
@@ 5. Further Work
In addition, we aim to reintegrate an articulatory–acoustic module into a neural model framework. This will allow DYNARTmo to serve as a component within a large-scale, neurobiologically plausible model of speech production and perception (Kröger,, 2023).

### v5

# DYNARTmo: A Dynamic Articulatory Model for Visualization of Speech Movement Patterns

@@ 1. Introduction
Articulatory models describe the stage of speech where neuromuscular activation patterns lead to movement patterns of speech articulators (i.e., lips, tongue, velum, lower jaw) with the goal of generating audible speech signals as part of speech communication processes. The motivation for developing articulatory models is manifold. (i) Models unfold the process of speech articulation and acoustic signal generation and thus help to understand the basic mechanisms of articulation and coarticulation. (ii) If coupled with an articulatory-acoustic module, an articulatory model provides insight into the mechanisms of acoustic signal generation at the vocal folds and within the vocal tract. (iii) Articulatory models may serve in the future as high-quality speech synthesizers, capable of generating different voices (i.e., different speakers), different emotional connotations, and more—beyond merely transmitting a spoken message. However, corpus-based synthesis systems currently produce the most natural speech quality, and many research questions remain to be addressed before high-quality articulatory synthesizers can be realized (Campbell2005; Kroger2022).
@@
At present, articulatory models are mainly used as research tools to study articulation and coarticulation or, when combined with articulatory-acoustic modules, to investigate aspects of acoustic signal generation, including sublaryngeal, laryngeal, and supralaryngeal aerodynamics (Birkholz2013; Kroger2022; Fan2024). Other models are used in speech therapy to visually present articulator positions ((Kroger2005)) or articulatory contact information (e.g., tongue–palate or labial contact), offering somatosensory feedback to help improve our imagination of articulation. Examples include electropalatography (EPG), which provides real‑time visualization of tongue–palate contact during speech (e.g. (Hardcastle1991; Nordberg2011)), and ultrasound visual biofeedback (U‑VBF), displaying dynamic tongue surface movements in therapy for speech sound disorders ((Sugden2019; Preston2017)). Moreover, articulatory models can be used to simulate speech learning or relearning processes in therapeutic contexts, or as part of second and third language learning. In that case, articulatory models need to be integrated into large-scale neurobiologically motivated neural models of language processing (Kroger2023).
@@
Many aspects of speech articulation can be represented in the midsagittal plane. Therefore, early articulatory models were typically two-dimensional (Henke1966; Mermelstein1973; Coker1976; Maeda1979; Heike1979). These models primarily aimed at generating high-quality synthetic speech. However, by the end of the 20th century, corpus-based synthesizers such as the Klatt synthesizer (Klatt1980) and newer systems like corpus-bases speech synthesizers (Campbell2005) had outperformed articulatory models in speech quality. Consequently, articulatory models—including articulatory-acoustic variants—became more focused on their use as research tools, particularly in studying speech acquisition, articulation, and coarticulation across languages and speaker groups. As a result, increasingly detailed models emerged that incorporate 3D spatial information, neuromuscular control schemes, and biomechanical modeling of muscles and articulatory tissues (Kroger2022).
@@
The model described in this paper, DYNARTmo (DYNamic ARTiculatory model), is a further development of the UK-DYNAMO (Universität Köln-DYNamic Articulatory MOdel) originally introduced by Heike (Heike1979). The UK-DYNAMO was inspired by Henke’s approach (Henke1966) and later used in a computer-based tool for speech training (Heike1989). Like many models of its time, UK-DYNAMO was a 2D system based on midsagittal contours of articulator surfaces, specifically German vowel and consonant contours from Wängler (Waengler1958). It employed temporal interpolation and a basic form of coarticulation as suggested by Henke.
@@
DYNARTmo was initially introduced by Kröger (Kroger1992), who incorporated the concept of articulatory underspecification to derive minimal rules for articulation and coarticulation. This approach was elaborated through two control concepts: a gestural control scheme (Kroger1993; KrogerEtAl1995) and a segmental approach (Kroger2001). Starting in the 2000s, DYNARTmo was updated with geometries derived from MRI data (KrogerEtAl2000; KrogerEtAl2004), offering a more realistic articulatory basis. Later, it was embedded into a large-scale neurobiological speech production model (KrogerEtAl2014), incorporating concepts such as primary and secondary articulators, muscle bundles, and neuromuscular activation patterns (KrogerBekolay2022).
@@ 2. DYNARTmo: Static Aspects
Table 1: Continuously scaled control parameters used in DYNARTmo, including affected articulators and their functional roles. Vocalic parameters define global vocal tract shaping; consonantal parameters govern localized constrictions. Additional parameters control velopharyngeal aperture, glottal configuration, and subglottal pressure. (Speech sounds are transcribed using SAMPA; see (Wells1997))
@@
In addition to the ten continuously scaled control parameters (see Table 1), the model includes six discrete parameters that define the place and manner of consonantal constrictions (see Table 2). The value range of the continuous parameters varies depending on their functional category: for vocalic parameters, values range from 1-1 to $1$, where 0 indicates the neutral articulator position. The same 1-1 to $1$ scale is used for velopharyngeal and glottal aperture, with $1$ indicating full opening, 0 indicating closure, and 1-1 representing tight closure. For parameters governing consonantal constrictions (i.e., labial, apical, and dorsal), as well as for vocal fold tension and lung pressure, the scale ranges from 0 to $1$. A value of 0 indicates no active constriction or minimum effort (i.e., the articulator remains in its current vocalic position), while a value of $1$ represents full articulatory closure or maximum effort.
@@
tongue tip manner full (vocalic), near (central groove), lateral
@@
The set of articulatory control parameters does not directly determine the positions of the model articulators. Rather, these parameters serve a functional role: they specify consonantal articulation in terms of the degree and location of vocal tract constrictions, and vocalic articulation in terms of canonical phonetic-linguistic categories such as high–low, front–back, and spread–rounded (cf.(IPA1999): International Phonetic Association).
The actual specification of consonantal articulator positions is carried out in a subsequent computational step, based on the values of all control parameters at a given time point. This step involves computing the vertical and horizontal positions of the tongue tip, tongue dorsum, and lips, as well as the vertical positions of the velum, lower jaw, and larynx (see Fig. 1). Conceptually, this transformation can be understood as the mapping of high-level vocal tract parameters—such as constriction degree and location for consonants, or vocal tract tube shape for vowels—onto concrete articulator geometries, akin to the task dynamics approach proposed by (SaltzmanMunhall1989).
@@ 3. Integration of DYNARTmo into a Web App for Speech Therapy and Language Learning
In addition to displaying the form and displacement of articulators from their rest positions in a sagittal plane view (see Fig. 1), the SpeechArticulationTrainer web application also provides a glottal view (transverse plane, superior perspective), and a palatal view (transverse plane, inferior perspective), as illustrated in Figure 2.
@@ 4. Discussion
One central goal of this study is to propose a simplified articulatory model that operates without the full task-dynamics formalism, while still reflecting key distinctions commonly found in more complex frameworks such as the Task Dynamics Approach ((SaltzmanMunhall1989)). Notably, our model implicitly includes a distinction between articulator space and tract variable space. However, this distinction is not strictly separated in our formulation. For instance, articulatory parameters such as tongue dorsum height or velum position are directly mapped to constriction properties, thereby bypassing an explicit intermediate representation of tract variables. This deliberate simplification serves to make the model computationally tractable and visually transparent, especially for educational or diagnostic applications. Nonetheless, the blending of articulator-based and functionally-defined parameters might lead to a certain conceptual ”blurring” that deserves clarification. Future refinements of the model may aim to more clearly disentangle these parameter spaces, or at least highlight their interplay more explicitly, in order to facilitate theoretical comparisons with established models of speech production.
@@ 5. Further Work
In addition, we aim to reintegrate an articulatory–acoustic module into a neural model framework. This will allow DYNARTmo to serve as a component within a large-scale, neurobiologically plausible model of speech production and perception (Kroger2023).